OpenStudy (anonymous):

Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this. -3x + y = 4 x + 3y = 22 Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this. -3x + y = 4 x + 3y = 22 @Mathematics

OpenStudy (saifoo.khan):

x = 1 y = 7

OpenStudy (anonymous):

I don't get how you got that. I've done it like three times and got zero for the first equation and it doesn't sound right

OpenStudy (anonymous):

I'll will show you how to solve it. Give me just a few minutes.

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

Solve one of the equations for either x or y. I chose to solve -3x + y = 4 for y. -3x + y = 4 -3x + 3x + y = 4 + 3x y = 4 + 3x Now you take the value of y which is y = 4 + 3x and substitute that into the other equation and solve for x, like so: x + 3y = 22 x + 3 (4 + 3x) = 22 x + 12 + 9x = 22 10x + 12 = 22 10x + 12 - 12 = 22 - 12 10x = 10 10x/10 = 10/10 x = 1 Now that you have the value for x you can solve for y in the other equation. -3x + y = 4 -3(1) + y = 4 -3 + y = 4 -3 + 3 + y = 4 + 3 y = 7 Now check the values of x and y by substituting them into both equations and seeing if they equal out. -3x + y = 4 -3(1) + 7 = 4 -3 + 7 = 4 4 = 4 x + 3y = 22 1 + 3 (7) = 22 1 + 21 = 22 22 = 22 The solution checks out. So your answer is x = 1 and y = 7

OpenStudy (anonymous):

Okay see i had the steps right I just didn't have the one instead i was using a zero and getting myself confused thank you so much for the help

OpenStudy (anonymous):

You are very welcome. :)